# Unsupervised Document and Template Clustering using Multimodal Embeddings

### v1

@@ 4.3.2 Experimental results
In the mixed-corpus experiment, the global hierarchy differs sharply from the template-level picture. Table 7 summarizes the results of document-level clustering, showing how effectively each model–algorithm pair separates pages by category. When the number of clusters is known, ColPali reaches perfect concordance with the ground truth (ARI = 1, NMI = 1, HS = CS = 1). The contrast in character sets, languages and syntactic markers across datasets is therefore leveraged to line up each document class with a distinct centroid. Purely textual SBERT follows closely (ARI = 0.989, NMI = 0.984) as vocabulary and orthography differ markedly between Brazilian civil IDs, Mexican death certificates, and English receipts. Token identity alone is therefore highly discriminative. The vision–text hybrid Donut also clusters the documents almost flawlessly (ARI = 0.944) despite seeing only the raster at inference time. The autoregressive coupling between local pixels and the textual decoder evidently transfers enough lexical awareness to the encoder representation.
@@
Table 7: Document-level clustering results on the six-class mixed corpus. Metrics are adjusted Rand index (ARI), normalized mutual information (NMI), homogeneity score (HS), completeness score (CS), V-measure (VM), silhouette score (SS), the number of predicted clusters (PC) – ground truth has 6 clusters – and the percentage of noise points returned by DBSCAN. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS VM SS PC % Noise
SBERT $k$-Means 0.9890 0.9837 0.9837 0.9837 0.9837 0.3964
SBERT DBSCAN 0.5929 0.7750 0.6921 0.8804 0.7750 0.3793 5 4.72
LayoutLMv1 $k$-Means 0.7094 0.8530 0.8248 0.8833 0.853 0.4763
LayoutLMv1 DBSCAN 0.6910 0.8238 0.8289 0.8187 0.8238 0.4703 7 7.27
LayoutLMv3 $k$-Means 0.7881 0.8419 0.8415 0.8422 0.8419 0.4568
LayoutLMv3 DBSCAN 0.5805 0.7205 0.7319 0.7094 0.7205 0.4522 8 4.32
DiT $k$-Means 0.3795 0.5025 0.4973 0.5079 0.5025 0.3231
DiT DBSCAN 0.0810 0.2647 0.1689 0.6115 0.2647 0.2278 3 13.19
Donut $k$-Means 0.9437 0.9471 0.9467 0.9475 0.9471 0.6253
Donut DBSCAN 0.8823 0.9126 0.9702 0.8614 0.9126 0.6110 9 6.55
ColPali $k$-Means 1 1 1 1 1 0.5581
ColPali DBSCAN 0.7337 0.8457 0.8479 0.8435 0.8457 0.5344 10 8.23
The multimodal but layout-oriented LayoutLM family occupies a middle ground. LayoutLMv1 reaches ARI = 0.709, while LayoutLMv3 delivers ARI = 0.788, benefitting from patch-level vision. The vision-only DiT trails far behind (ARI = 0.380). The six classes share many graphic primitives — official crests, ruled boxes, QR codes — so purely visual tokens provide insufficient separability without textual anchors.
When the oracle assumption is removed, DBSCAN exposes qualitatively different failure modes. Donut retains the lead (ARI = 0.882) and labels only 6.6% of pages as noise; its clusters are geometrically tight (HS = 0.97) but slightly over-segmented (CS = 0.86), producing seven clusters, two of which split the BID sub-set along subtle typography variants. ColPali still performs strongly (ARI = 0.734) but expels 8.2% of the pages. SBERT (ARI = 0.593) and LayoutLMv1 (ARI = 0.691) mainly suffer from under-segmentation: five clusters are returned, with FATURA and SROIE often merged because both contain Latin corporate headers and dense monetary amounts. DiT once again illustrate how vision-only encoders are brittle when neighbourhood density must be chosen globally: the DiT manifold collapses to three macro-regions, forcing DBSCAN to discard 13% of the documents as outliers. Finally, LayoutLMv3 presents good performance as expected (V-measure = 0.7205), although inferior to Donut (V-measure = 0.9126) and ColPali (V-measure = 0.8457).
@@
Figure 4: Two–dimensional t-SNE projections of the mixed-corpus dataset using the six embedding models. Colors denote document categories; each plot therefore illustrates how well the corresponding representation separates distinct classes in the document-level clustering setting.
The qualitative structure revealed by the t-SNE projections in Figure 4 mirrors the quantitative rankings reported in Table 7. Embeddings produced by ColPali and SBERT give rise to compact, approximately isotropic clusters that are cleanly isolated along the manifold, visually corroborating their highest ARI, NMI and V-measure scores. LayoutLMv1 and Donut retain the same coarse class separation, but their clusters display slight elongation and occasional boundary fraying, consistent with the small performance gap they show relative to the top tier. In the LayoutLMv3 projection, several categories appear fragmented into multiple lobes and radial filaments, anticipating the moderate completeness and homogeneity reported for this model. Finally, DiT produces largely amorphous clouds with extensive class overlap, a pattern that echoes its markedly lower document-level clustering scores. The geometric organization observed in two dimensions therefore reinforces the conclusion that visual-textual fusion and text-centric representations dominate document-level discrimination, whereas the purely visual DiT encoder remains ill-suited to this task.
@@ 4.4.1 Size of the encoder architecture
Table 8: Ablation study: LayoutLMv1 base versus LayoutLMv1 large on mixed-corpus document-level clustering.
Model Clustering Alg. ARI NMI HS CS VM SS PC % Noise
LayoutLMv1 base $k$-Means 0.7094 0.8530 0.8248 0.8833 0.8530 0.4763
LayoutLMv1 base DBSCAN 0.6910 0.8238 0.8289 0.8187 0.8238 0.4703 7 7.27
LayoutLMv1 large $k$-Means 0.8692 0.8908 0.8877 0.8908 0.8908 0.4301
LayoutLMv1 large DBSCAN 0.5951 0.7845 0.7018 0.8893 0.7845 0.4701 5 5.12
Under oracle $k$-Means the larger network provides the expected margin gain: adjusted Rand index rises from 0.71 to 0.87, normalized mutual information increases from 0.85 to 0.89 and both homogeneity and completeness surpass 0.88, confirming that additional depth sharpens the global separation among dataset categories. However, when the number of clusters is unknown, the results are reversed. On DBSCAN every metric except V-measure is worse for the large model. The base representation yields 0.69 ARI, 0.82 NMI, 0.83 homogeneity, 0.82 completeness, a silhouette of 0.48, seven clusters discovered, and 7.3% noise. Scaling up collapses two classes into shared density basins, leaving five clusters, driving ARI down to 0.60, NMI to 0.78, and homogeneity to 0.70, while reducing noise only marginally to 5.1%. V-measure alone shows a moderate decrease (from 0.82 to 0.78 because completeness improved by 9 points), but the class-merging error dominates all other criteria.
These results show that, while larger models improve the overall separation between classes, they also smooth out the fine details in the data that density-based methods such as DBSCAN rely on. The divergence between V-measure and the other metrics underscores the importance of joint evaluation. Practically, the results caution that scaling a multimodal transformer is beneficial when the cluster cardinality is known, but may hinder open-world discovery unless multi-scale pooling preserves local density structure.
@@ 4.4.2 Mean pooling versus [CLS] token
Table 9: Ablation study: mean pooling versus [CLS] token on mixed-corpus document-level clustering.
Approach Clustering Alg. ARI NMI HS CS VM SS PC % Noise
Mean pooling $k$-Means 0.7094 0.8530 0.8248 0.8833 0.8530 0.4763
Mean pooling DBSCAN 0.6910 0.8238 0.8289 0.8187 0.8238 0.4703 7 7.27
[CLS] $k$-Means 0.5513 0.6515 0.6475 0.6556 0.6515 0.5272
[CLS] DBSCAN 0.4658 0.6332 0.6308 0.6357 0.6332 0.4932 7 13.35
When the number of clusters $k$ is supplied, mean pooling reaches 0.7094 ARI and 0.853 NMI, whereas the [CLS] vector attains only 0.5513 ARI and 0.6515 NMI. Homogeneity and completeness decrease from 0.8248 and 0.8833 to 0.6475 and 0.6556, respectively. With DBSCAN the gap widens: mean pooling preserves 0.691 ARI and 0.8238 NMI, while the [CLS] variant falls to 0.4658 ARI, 0.6332 NMI. V-measure also decreases significantly from 0.8238 to 0.6332.

### v3

@@ 4.4.2 Experimental Results
In the mixed-corpus experiment, the global hierarchy differs sharply from the template-level picture. Table 7 summarizes the results of document-level clustering, showing how effectively each model–algorithm pair separates pages by category.
$k$-Means.
With the true cardinality supplied, ColPali reaches perfect concordance with the ground truth, achieving ARI = 1, NMI = 1, HS = 1, CS = 1, SS = 0.5581, indicating that its manifold is linearly separable at the class scale. The SBERT baseline is nearly perfect as well, ARI = 0.9890, NMI = 0.9837, HS = 0.9837, CS = 0.9837, SS = 0.3964, showing that token identity and language cues alone suffice to distinguish these six categories when $k$ is known. Among multimodal encoders, Donut performs strongly (ARI = 0.9437, NMI = 0.9471, HS = 0.9467, CS = 0.9475, SS = 0.6253), with the highest $k$-Means silhouette among the top performers, reflecting compact and well-separated clusters. Gemma3 clusters well under the oracle with ARI = 0.7853 and NMI = 0.8427, and its silhouette of 0.6596 is the highest among all $k$-Means runs, suggesting especially tight intra-cluster geometry despite moderate ARI. LayoutLMv3 presents similar external agreement with ARI = 0.7881 and NMI = 0.8419. LayoutLMv1 trails these two with ARI = 0.7094, NMI = 0.8530, HS = 0.8248, CS = 0.8833, SS = 0.4763; its high completeness relative to homogeneity hints at occasional under-segmentation across visually similar categories. InternVL3 performs less well overall (ARI = 0.5195, NMI = 0.6631), yet it achieves a notably high silhouette score (SS = 0.6631), suggesting that its clusters are internally coherent despite limited alignment with the ground-truth labels. Purely visual DiT ranks last under oracle $k$ (HS = 0.4973, CS = 0.5079), consistent with the notion that document categories are largely driven by lexical and symbolic differences that vision-only embeddings struggle to separate.
BIRCH.
With the same advantage of the true $k$, BIRCH reproduces the $k$-Means hierarchy with minor differences. ColPali again achieves perfect scores on all external metrics. SBERT slightly improves to ARI = 0.9909. Donut remains strong with NMI = 0.9524, HS = 0.9514 and CS = 0.9535, while Gemma3 edges up relative to its $k$-Means run, obtaining ARI = 0.8154 and NMI = 0.8725. LayoutLMv1 improves modestly, while LayoutLMv3 is worse under BIRCH than under $k$-Means, suggesting that its manifold responds more favorably to centroid partitioning than to BIRCH’s incremental clustering. InternVL3 and DiT increase slightly relative to $k$-Means, but both remain far from the top tier. Under oracle conditions, the overall picture is therefore stable: language-dominated or highly aligned vision–language embeddings cluster categories almost perfectly; visual-only signals are insufficient; and BIRCH gives broadly comparable results to $k$-Means, with small model-dependent shifts.
@@
Table 7: Document-level clustering results on the six-class mixed corpus. Metrics are adjusted rand index (ARI), normalized mutual information (NMI), homogeneity score (HS), completeness score (CS), silhouette score ($S$), the number of predicted clusters (PC) and the percentage of noise points returned by DBSCAN. Ground truth has 6 clusters. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS $S$ PC % Noise
SBERT $k$-Means 0.9890 0.9837 0.9837 0.9837 0.3964
SBERT DBSCAN 0.5929 0.7750 0.6921 0.8804 0.3793 5 4.72
SBERT HDBSCAN + $k$-NN 0.8185 0.9152 0.8574 0.9813 0.4107 5
SBERT BIRCH 0.9909 0.9872 0.9872 0.9872 0.3930
LayoutLMv1 $k$-Means 0.7094 0.8530 0.8248 0.8833 0.4763
LayoutLMv1 DBSCAN 0.6910 0.8238 0.8289 0.8187 0.4703 7 7.27
LayoutLMv1 HDBSCAN + $k$-NN 0.7327 0.8534 0.8287 0.8796 0.4747 6
LayoutLMv1 BIRCH 0.7272 0.8662 0.8387 0.8955 0.4767
LayoutLMv3 $k$-Means 0.7881 0.8419 0.8415 0.8422 0.4568
LayoutLMv3 DBSCAN 0.5805 0.7205 0.7319 0.7094 0.4522 8 4.32
LayoutLMv3 HDBSCAN + $k$-NN 0.4495 0.6455 0.8055 0.5385 0.4131 25
LayoutLMv3 BIRCH 0.5597 0.6884 0.6680 0.7101 0.5138
DiT $k$-Means 0.3795 0.5025 0.4973 0.5079 0.3231
DiT DBSCAN 0.0810 0.2647 0.1689 0.6115 0.2278 3 13.19
DiT HDBSCAN + $k$-NN 0.0483 0.2551 0.1793 0.4415 0.1048 11
DiT BIRCH 0.4710 0.6091 0.5832 0.6374 0.3260
Donut $k$-Means 0.9437 0.9471 0.9467 0.9475 0.6253
Donut DBSCAN 0.8823 0.9126 0.9702 0.8614 0.6110 9 6.55
Donut HDBSCAN + $k$-NN 0.8503 0.8790 0.9815 0.7958 0.5352 15
Donut BIRCH 0.9421 0.9524 0.9514 0.9535 0.6195
ColPali $k$-Means 1 1 1 1 0.5581
ColPali DBSCAN 0.7337 0.8457 0.8479 0.8435 0.5344 10 8.23
ColPali HDBSCAN + $k$-NN 0.7043 0.8583 0.7518 1 0.6500 4
ColPali BIRCH 1 1 1 1 0.5581
InternVL3 $k$-Means 0.5195 0.6631 0.6309 0.6987 0.6631
InternVL3 DBSCAN 0.6110 0.7172 0.7892 0.6573 0.5934 12 2.64
InternVL3 HDBSCAN + $k$-NN 0.6364 0.8025 0.9972 0.6714 0.5204 19
InternVL3 BIRCH 0.5294 0.6513 0.6365 0.6668 0.5787
Gemma3 $k$-Means 0.7853 0.8427 0.8419 0.8435 0.6596
Gemma3 DBSCAN 0.6725 0.7949 0.9354 0.6911 0.4371 17 12.15
Gemma3 HDBSCAN + $k$-NN 0.7362 0.8345 0.9945 0.7188 0.5937 17
Gemma3 BIRCH 0.8154 0.8725 0.8719 0.8730 0.6556
DBSCAN.
Removing the oracle constraint changes the landscape. Donut now leads the field with ARI = 0.8823, NMI = 0.9126, HS = 0.9702, CS = 0.8614, SS = 0.6110, PC = 9, noise = 6.55%, i.e., slightly over-segmenting (nine clusters versus six) while maintaining very high homogeneity, a signature of tight density basins that occasionally split a class. ColPali continues to perform well, once more exhibiting slight over-segmentation. Gemma3 achieves ARI = 0.6725, NMI = 0.7949, HS = 0.9354, CS = 0.6911, SS = 0.4371, PC = 17, noise = 0.1215. The sharp divergence between its very high homogeneity and lower completeness indicates over-segmentation into many pure fragments, which aligns with the predicted 17 clusters and non-trivial noise. InternVL3 follows with ARI = 0.6110, NMI = 0.7172, HS = 0.7892, CS = 0.6573, SS = 0.5934, PC = 12, noise = 2.64%, exhibiting the same over-fragmentation pattern but with less discarded documents. LayoutLMv1 preserves good balance with ARI = 0.6910 and NMI = 0.8238, HS = 0.8289, CS = 0.8187, SS = 0.4703, PC = 7, noise = 0.0727, and LayoutLMv3 is somewhat lower with ARI = 0.5805 and NMI = 0.7205. SBERT drops markedly compared to oracle $k$, achieving ARI = 0.5929, NMI = 0.7750, HS = 0.6921, CS = 0.8804, SS = 0.3793, PC = 5, noise = 4.72%, showing under-segmentation. DiT degrades most (ARI = 0.0810, NMI = 0.2647), highlighting the brittleness of purely visual tokens when a single global density scale must explain heterogeneous categories.
HDBSCAN + $k$-NN.
The hybrid density-then-assignment scheme eliminates outliers by design and exposes a clear split between under-segmentation and over-segmentation behaviors depending on the embedding. SBERT returns ARI = 0.8185, NMI = 0.9152, HS = 0.8574, CS = 0.9813, SS = 0.4107, PC = 5. The near-perfect completeness with fewer than six clusters indicates merging of classes, consistent with its DBSCAN failure mode. LayoutLMv1 remains balanced with ARI = 0.7327 and NMI = 0.8534, effectively recovering the oracle cardinality. ColPali collapses to four clusters and markedly lower homogeneity, a canonical signature of under-segmentation. In contrast, Donut over-segments (ARI = 0.8503, NMI = 0.8790, HS = 0.9815, CS = 0.7958, SS = 0.5352, PC = 15), creating many very pure clusters that fragment classes. The same pattern is pronounced for the two large vision-language models. Gemma3 yields HS = 0.9945 and CS = 0.7188, and InternVL3 obtains HS = 0.9972 and CS = 0.6714, both showing extreme over-segmentation with near-unity homogeneity but depressed completeness. LayoutLMv3 also over-segments sharply, while DiT performs the worst (ARI = 0.0483, NMI = 0.2551), again reflecting the limitations of visual tokens without textual anchors in this multi-corpus setting.
Embedding Geometry via t-SNE.
The qualitative structure revealed by the t-SNE projections in Figure 4 mirrors the quantitative rankings reported in Table 7. Embeddings produced by ColPali and SBERT give rise to compact, approximately isotropic clusters that are cleanly isolated along the manifold, visually corroborating their highest ARI and NMI scores. LayoutLMv1 and Donut retain the same coarse class separation, but their clusters display slight elongation and occasional boundary fraying, consistent with the small performance gap they show relative to the top tier. In the LayoutLMv3 projection, several categories appear fragmented into multiple lobes and radial filaments, anticipating the moderate completeness and homogeneity reported for this model. DiT produces largely amorphous clouds with extensive class overlap, a pattern that echoes its markedly lower document-level clustering scores.
The projection derived from Gemma3 reveals tightly packed and highly localized clusters, with well-separated class regions and minimal inter-cluster interference. Despite some variation in intra-cluster density, the overall configuration is spatially organized and reflects the model’s strong internal cohesion observed in its high silhouette scores, even in settings where it over-segments. In contrast, InternVL3 exhibits more complex geometry: while many clusters remain distinct, the configuration is less uniform, with several class fragments positioned at varying distances and some evidence of discontinuities. This matches its tendency to form compact, low-noise clusters that only partially align with ground-truth labels, as captured by its moderate ARI and NMI values. Together, the t-SNE projections of these newer models confirm that they generate internally coherent structures, though their global label alignment varies according to the clustering objective.
The geometric organization observed in two dimensions therefore reinforces the conclusion that visual-textual fusion and text-centric representations dominate document-level discrimination, whereas the purely visual DiT encoder remains ill-suited to this task.
@@
(g) Gemma3
(h) InternVL3
Figure 4: Two–dimensional t-SNE projections of the mixed-corpus dataset using the eight embedding models. Colors denote document categories; each plot therefore illustrates how well the corresponding representation separates distinct classes in the document-level clustering setting.
Interpreting Document-Level Clustering Dynamics.
The document-level corpus exposes three clear patterns. First, when provided with the oracle number of clusters, $k$-Means and BIRCH make the task almost trivial for language‑centric or well‑aligned vision–language embeddings: ColPali is exact (all external metrics = 1), SBERT is nearly perfect, and Donut is close behind.
Second, removing the oracle reveals modality‑dependent failure modes. DBSCAN cleanly separates under‑segmentation from over‑segmentation: SBERT and ColPali tend to merge classes (five and ten clusters, respectively), whereas Donut, Gemma3, and InternVL3 fragment classes (roughly 9, 17, and 12 clusters), often with high homogeneity but reduced completeness. Noise remains moderate for top models (e.g., Donut = 6.6%, SBERT = 4.7%) and higher for Gemma3 and DiT (12% and 13%, respectively). Even when fragmented, Donut and InternVL3 retain high silhouettes, indicating internally coherent pieces.
Third, HDBSCAN + $k$-NN further clarifies the trade‑off between merging and splitting. ColPali under‑segments, while Donut, Gemma3, and InternVL3 produce many very pure islands. LayoutLMv1 is a notable exception, recovering the true cardinality with balanced homogeneity/completeness. The absence of noise highlights that assigning every point typically trades off purity through over-segmentation or stability through class merging.
From a design perspective, these observations argue that document-level clustering benefits most from embeddings that preserve token identity and language, with vision and layout contributing decisively when typography and structure discriminate classes but not when categories differ primarily by content. DiT’s weaker non‑oracle results reinforce this point.
In practical applications, the gap between oracle and non-oracle settings is critical. Since the true number of clusters is rarely known in advance, strong multimodal embeddings must be paired with mechanisms such as robust $k$-estimation, adaptive density calibration, or post hoc consolidation strategies. Ultimately, language-centric or well-aligned vision and language models offer the most dependable foundation for unsupervised category discovery, but their effectiveness depends on careful alignment between the embedding space and the clustering strategy used.
@@ 4.5.1 Size of the Encoder Architecture
Table 8: Ablation study: LayoutLMv1 base versus LayoutLMv1 large on mixed-corpus document-level clustering. Metrics are adjusted rand index (ARI), normalized mutual information (NMI), homogeneity score (HS), completeness score (CS), silhouette score ($S$), the number of predicted clusters (PC) and the percentage of noise points returned by DBSCAN. Ground truth has 6 clusters. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS $S$ PC % Noise
LayoutLMv1 base $k$-Means 0.7094 0.8530 0.8248 0.8833 0.4763
LayoutLMv1 base DBSCAN 0.6910 0.8238 0.8289 0.8187 0.4703 7 7.27
LayoutLMv1 large $k$-Means 0.8692 0.8908 0.8877 0.8908 0.4301
LayoutLMv1 large DBSCAN 0.5951 0.7845 0.7018 0.8893 0.4701 5 5.12
Under oracle $k$-Means the larger network provides the expected margin gain: adjusted Rand index rises from 0.71 to 0.87, normalized mutual information increases from 0.85 to 0.89 and both homogeneity and completeness surpass 0.88, confirming that additional depth sharpens the global separation among dataset categories. However, when the number of clusters is unknown, the results are reversed. On DBSCAN every metric except completeness is worse for the large model. The base representation yields 0.69 ARI, 0.82 NMI, 0.83 homogeneity, 0.82 completeness, a silhouette of 0.48, seven clusters discovered, and 7.3% noise. Scaling up collapses two classes into shared density basins, leaving five clusters, driving ARI down to 0.60, NMI to 0.78, and homogeneity to 0.70, while reducing noise only marginally to 5.1%.
These results show that, while larger models improve the overall separation between classes, they also smooth out the fine details in the data that density-based methods such as DBSCAN rely on. The divergence between completeness and the other metrics underscores the importance of joint evaluation. Practically, the results caution that scaling a multimodal transformer is beneficial when the cluster cardinality is known, but may hinder open-world discovery unless multi-scale pooling preserves local density structure.
@@ 4.5.2 Mean Pooling Versus [CLS] Token
Table 9: Ablation study: mean pooling versus [CLS] token on mixed-corpus document-level clustering. Metrics are adjusted rand index (ARI), normalized mutual information (NMI), homogeneity score (HS), completeness score (CS), silhouette score ($S$), the number of predicted clusters (PC) and the percentage of noise points returned by DBSCAN. Ground truth has 6 clusters. The top-performing scores are shown in boldface.
Approach Clustering Alg. ARI NMI HS CS $S$ PC % Noise
Mean pooling $k$-Means 0.7094 0.8530 0.8248 0.8833 0.4763
Mean pooling DBSCAN 0.6910 0.8238 0.8289 0.8187 0.4703 7 7.27
[CLS] $k$-Means 0.5513 0.6515 0.6475 0.6556 0.5272
[CLS] DBSCAN 0.4658 0.6332 0.6308 0.6357 0.4932 7 13.35
When the number of clusters $k$ is supplied, mean pooling reaches 0.7094 ARI and 0.853 NMI, whereas the [CLS] vector attains only 0.5513 ARI and 0.6515 NMI. Homogeneity and completeness decrease from 0.8248 and 0.8833 to 0.6475 and 0.6556, respectively. With DBSCAN the gap widens: mean pooling preserves 0.691 ARI and 0.8238 NMI, while the [CLS] variant falls to 0.4658 ARI, 0.6332 NMI.